# Should you use LLMs to simulate opinions? Quality checks for early-stage deliberation

### v3

@@ Abstract
The emergent capabilities of large language models (LLMs) have prompted interest in using them as surrogates for human subjects in opinion surveys. However, prior evaluations of LLM-based opinion simulation have relied heavily on costly, domain-specific survey data, and mixed empirical results leave their reliability in question. To enable cost-effective, early-stage evaluation, we introduce a quality control assessment designed to test the viability of LLM-simulated opinions on Likert-scale tasks without requiring large-scale human data for validation. This assessment comprises two key tests: logical consistency and alignment with stakeholder expectations, offering a low-cost, domain-adaptable validation tool. We apply our quality control assessment to an opinion simulation task relevant to AI-assisted content moderation and fact-checking workflows—a socially impactful use case—and evaluate seven LLMs using a baseline prompt engineering method (backstory prompting), as well as fine-tuning and in-context learning variants. None of the models or methods pass the full assessment, revealing several failure modes. We conclude with a discussion of the risk management implications and release TopicMisinfo, a benchmark dataset with paired human and LLM annotations simulated by various models and approaches, to support future research.
@@ Introduction
Large language models (LLMs) have demonstrated significant appeal and versatility across a wide range of tasks. While the primary training objective of the base model is to predict the next most likely token, training on vast and diverse datasets, combined with increasingly complex system architectures, has resulted in the “emergence” of capabilities that were not explicitly anticipated during development (Wei et al. 2022). For instance, LLMs have been shown to learn new tasks through in-context learning (Brown et al. 2020), perform coding tasks with remarkable accuracy (Zheng et al. 2024), and have displayed human-like capabilities across different tasks (Kosinski 2024; Strachan et al. 2024).
@@
These emerging capabilities have led researchers to investigate more unconventional uses of LLMs, including automating opinion surveys by simulating the viewpoints of specific demographic or ideological groups (Qu and Wang 2024). The possibility that LLMs may successfully simulate demographic viewpoints stems from (i) evidence that LLMs can be prompted to replicate responses from human subject studies in psychology and other social sciences (Aher, Arriaga, and Kalai 2023), (ii) training data that likely encompass a broad spectrum of perspectives and opinions (Miranda et al. 2024), and (iii) prompting techniques that lead LLMs to produce seemingly coherent, opinionated responses (Wright et al. 2024). Automating such surveys could have broad applications in marketing (Sarstedt et al. 2024), content moderation (Fröhling, Demartini, and Assenmacher 2024), policymaking, and public relations (Sanders, Ulinich, and Schneier 2023). LLMs may be especially valuable for pilot studies (Sarstedt et al. 2024; Rothschild et al. 2024), sampling hard-to-reach populations (Jansen, gyo Jung, and Salminen 2023), and labeling data that may be psychologically harmful for human annotators (Wang, Morgenstern, and Dickerson 2025).
@@
However, existing evidence is inconclusive on the extent to which LLM-based methods can accurately simulate human opinions. Some studies show that including a demographic “backstory” in a prompt (Argyle et al. 2023; Bui et al. 2025; Jiang, Wei, and Zhang 2024) (i.e., “backstory” prompting), fine‑tuning (Namikoshi et al. 2024), and in‑context learning (Karanjai et al. 2025) are approaches that can successfully approximate group‑level opinions, while others document systematic failures of these same approaches—especially when simulating the opinions of minority or non‑Western groups (Santurkar et al. 2023; Sun et al. 2023; Qu and Wang 2024; Mingmeng, He, and Trotta 2024; Orlikowski et al. 2025). Additional work finds that backstory prompting approaches—the most prevalent approach to simulating opinions in prior research—flatten within-group variance of opinions (Wang, Morgenstern, and Dickerson 2025; Bisbee et al. 2024; Mingmeng, He, and Trotta 2024) and fails to generalize across topics (Sanders, Ulinich, and Schneier 2023; Lee et al. 2024). See Appendix 1 for a literature review table that summarizes prior works methodologies, respective application domains, and findings.
@@
The mixed results in the empirical literature suggest that the reliability of LLM-simulated opinions cannot be assumed across domains and instead requires rigorous validation for each new use case. Importantly, however, current validation methods rely on gathering high-quality human-labeled data—the cost of which can rival that of traditional surveys. For early‑stage assessments, this upfront expense is likely to outweigh uncertain benefits, significantly hindering research and development efforts. The existence of applications with broad consensus about the potential desirability of simulating opinions, such as when human survey subjects would be exposed to harmful content (Wang, Morgenstern, and Dickerson 2025) or trade secrets (Wang, Zhang, and Zhang 2024), motivates the need for validation techniques that lower the cost of early-stage exploration of LLM usage to simulate opinions.
@@
    Logical Consistency—whether models reliably position “average” opinions as a convex combination of group-level opinions and/or whether these average opinions are sampled from a stable reference population.
@@
The design of the proposed quality control assessment is grounded in the fact that some of the challenges involved in validating LLM-simulated opinions are similar to those involved in validating subjective human annotations. Researchers have developed various quality controls to assess the reliability of human data and annotator attentiveness (Artstein and Poesio 2008; Sap et al. 2022; Mostafazadeh Davani, Díaz, and Prabhakaran 2022; Lease 2011). Our proposed tests draw from this literature, adapting it to be suitable for assessing LLMs, with a focus on Likert-scale opinions. In the discussion section, we provide recommendations for integrating these tests into organizational risk management practices.
The value of these tests compared to those proposed in previous work (Aher, Arriaga, and Kalai 2023; Argyle et al. 2023; Sun et al. 2023) is that they can be performed prior to expensive large-scale human validation effort, fostering well-informed deliberation on the suitability of LLM opinion simulation in a particular context.
@@
To illustrate the utility of our quality checks, we develop a domain-specific testbed in content moderation, focusing on the prioritization of potentially harmful misinformation for fact-checker review. Because prioritization is opinion-driven, resource-intensive, and exposes annotators to psychologically taxing content (Liu, Gwizdka, and Lease 2024), it provides a compelling, socially impactful context for evaluating simulated opinions.
An effective testbed for group-level opinion simulation should include topics with both expected agreement and disagreement amongst groups. Gender offers strong prior expectations based on a significant body of public opinion research (Huddy, Cassese, and Lizotte 2008; Lizotte 2020). We focus on the prioritization criterion of identifying “potential harm to specific groups,” as previous research highlights this dimension of prioritization as one along which gender differences are particularly pronounced  (Huddy, Cassese, and Lizotte 2008; Lizotte 2020). [^1] While not the only criteria relevant for prioritization, research has noted that this is an important axis of consideration used by fact-checkers (Sehat et al. 2024; Liu, Gwizdka, and Lease 2024).
We built TopicMisinfo, a 160‑claim dataset spanning topics that can be expected to elicit gender disagreement and consensus. The selection of these topics is anchored in findings from the American National Election Studies (ANES) (American National Election Studies 2021), which help us capture a diversity in topics that are likely to provoke varying degrees of gendered disagreement. We also included “Gold” comprehension checks, which are claims that are harmless, and are either obviously true (e.g., “A circle is round.”) or obviously false (e.g., “The tallest tree on Earth touches Mars.”), with a very strong prior on gender-level consensus of opinion. We also collect a small sample ($\approx 1600$ annotations) of task-specific human opinion data. Dataset composition, example claims, and human annotation statistics are detailed in the Appendix 2-3; the TopicMisinfo dataset is publicly available to support future research. [^2] Link removed for peer review
@@
Applying the proposed quality checks, we evaluate simulated opinions from several commercially available LLMs: GPT-3.5-Turbo (checkpoints 06-13-23, 11-06-23, 01-25-24), GPT-4, LLaMA 3, Titan-Text-Premier, and Mistral-Large—referred to throughout as GPT3.5a, GPT3.5b, GPT3.5c, GPT-4, LLaMA3, Titan, and Mistral. We elicit group-level responses via backstory prompts (Argyle et al. 2023; Sun et al. 2023) that instruct models to adopt personas of men or women (see Appendix 4 for prompt templates). We also benchmark two other approaches to simulating opinions suggested in the literature: fine-tuning on small-sample human survey data (Namikoshi et al. 2024); and in-context learning with human-labeled examples (Karanjai et al. 2025). Importantly, our findings show that while fine-tuning and in-context learning generally improve alignment with stakeholder expectations, they do not substantially outperform backstory prompting on logical consistency. These results indicate that all evaluated models and prompting approaches fail the proposed quality control assessment for the task considered, suggesting that current methods face limitations in their ability to produce stable, internally coherent simulations of human opinions.
@@
In this section, we introduce two quality checks designed for early-stage evaluation of LLM-simulated opinions. For each check, we detail the methodology and analyze how a range of LLMs perform across gender conditions using the “backstory prompting” approach (Argyle et al. 2023) on claims drawn from the TopicMisinfo dataset, focusing on a key opinion task relevant to fact-checking and content moderation workflows: assessing the potential harm of claims to specific groups. The code and reproducible pipeline are publicly available. [^3] Link removed for peer review
@@
Testing for logical consistency is a foundational tool for evaluating the quality of data from human annotators (Aruguete et al. 2019) and from data aggregation methods such as majority voting (Mostafazadeh Davani, Díaz, and Prabhakaran 2022). In this work, we adapt this principle to assess the internal coherence of LLM-simulated opinions across demographic groups.
@@
For each claim $c_{i}$, the model is prompted $n_{i}$ times under each condition (men, women, average) using a temperature setting of $\tau=0.5$, generating distributions of Likert-scale responses. A logically consistent average opinion should be expressible as a convex combination of the group-level opinions it seeks to summarize. In the context of gender (assuming only two: men and women) [^4] While we assume a binary gender, Appendices 5a and 5b contains pseudo-code for generalizing this methodology for $\geq 2$ groups, while Appendix 5c includes the detail of a replication study including non-binary genders simulated with GPT3.5c., that would entail for each claim $c_{i}$, the model’s average predicted mean $\mu_{\text{avg}}(c_{i})$ should satisfy:
@@
We reject $H_{0}$ at level $\alpha$ if $\hat{p}(p_{0})<\alpha$, concluding that the model fails to produce average-conditioned opinions that lie within the convex hull of group-level distributions at an acceptable rate. We test acceptable rate values $p_{0}\in[0.7,0.8,0.9,1.0]$. See Appendix 5a for pseudocode that generalizes this test for $\geq 2$ groups.
@@
From this mixture, we resample synthetic average-conditioned responses, compute $\hat{q}^{(b)}_{c_{i}}$, and aggregate across claims: $\hat{L}^{(b)}_{\omega}(q_{0})=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\left|\hat{q}_{c_{i}}^{(b)}-q_{0}\right|$. The $p$-value is $\Pr\left(\hat{L}^{(b)}_{\omega}(q_{0})>\hat{L}_{\omega}(q_{0})\right)$. This test is repeated for values of $q_{0}$ in $[0,1]$ at intervals of 0.05, applying a Bonferroni correction ($\alpha^{*}=0.0025$) to account for multiple comparisons. See Appendix 5b for pseudocode that generalizes this test for $\geq 2$ groups.
@@ Results.
Figure 1 reports feasible consistency thresholds ($p_{0}$) for the weak test and valid $q_{0}$ intervals for the strong test, indicating alignment with a consistent reference population. We highlight results for four socially divisive topics where group-level disagreement is well-documented. Given the lack of a gold standard for prompting average-conditioned or population-level perspectives, we evaluate consistency across four alternative prompts. See Appendix 4 for details.
  • GPT-3.5c: For threshold $p_{0}=0.7$, passes the weak test across all topics with Base Prompt 4, indicating basic geometric plausibility. For the strong test, it only passes the Abortion topic, with feasible mixture weights $q_{0}\in[0,0.55]$ or $[0,0.85]$ depending on prompt. Fails to produce a consistent average-conditioned reference population for Black Americans, Illegal Immigration, and LGBTQ.
  • LLaMA-3: Two prompts pass the weak test at $p_{0}=0.7$ for Illegal Immigration and all four prompts pass for LGBTQ, but fails on other topics. For the strong test, exhibits moderate success: feasible $q_{0}\in[0,0.5]$ for Black Americans across prompts, and up-samples opinions of men for Illegal Immigration ($q_{0}\in[0.85,1]$). Fails on Abortion and LGBTQ.
@@
  • Titan: Two prompts pass the weak test at $p_{0}\in[0.7,0.8]$ for Black Americans and one prompt passes for LGBTQ at $p_{0}=0.7$. On the strong test, Prompt 4 yields success on Black Americans with full-range plausibility ($q_{0}\in[0,1]$). Prompts 1 and 2 yield plausible $q_{0}\ \in[0,1]$ for LGBTQ, but it fails on both Abortion and Illegal Immigration.
@@
Assessing alignment with common sense or widespread expectations is standard for evaluating human annotations, e.g., via attention checks (Lease 2011). The same underlying idea can be leveraged when assessing the quality of LLM-generated annotations. In our setting, for instance, stakeholders—e.g., end-users or developers—can have clear intuitions about cases where group-level differences should or should not arise. Checking for misalignment on these cases can provide a powerful lens for evaluating the quality of LLM-generated annotations. If neglected, such misalignment can bias downstream tasks, leading to skewed decisions or misrepresentation of group perspectives. In high-stakes settings like claim prioritization for fact-checking, such errors carry serious ethical consequences.
@@
Table 2: Combined Results for Multiple Models
Topic gpt-35-23-0613 gpt-35-23-1106 gpt-35-24-0125 gpt-4 llama3 mistral-large titan-text-premier
$\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val
@@
Abortion 1.65 * 0.28 0.60 -1.70 *** -1.65 *** -1.24 ** -1.16 *
Black Americans 1.72 * 0.35 0.39 -1.28 *** -1.09 *** -1.92 * -0.97 **
Illegal Immigration 0.92 ** -1.06 * -0.61 -1.42 ** -0.80 * -0.43 -0.71 *
LGBTQ 0.66 -1.37 * -0.37 -0.92 * -1.08 ** -0.99 ** -1.10 **
@@
Entertainment 1.59 * 1.05 * 1.18 * 0.35 ** 0.41 ** 0.58 * 0.48 *
Gold 0.43 0.34 0.71 ** -0.01 -0.30 0.02 0.32
HealthScience 1.86 * 1.07 * 0.65 -0.19 0.04 0.79 0.56
Sports -0.01 -0.06 0.21 -0.26 -0.26 -0.26 -0.16
USA 1.68 *** 1.12 *** 0.63 ** -0.61 *** -0.21 -0.56 * 0.16
WeatherClimate 1.06 * 1.27 * 0.97 * -0.39 -0.10 -0.68 0.10
@@
Table 3: Quality Check 2b Statistical Results. GPT-3.5-turbo models tend to significantly exaggerate gender differences on topics humans found less divisive, while GPT-4, Llama3, mistral-large, and Titan tended to significantly erode differences on topics humans found more divisive.
@@
a standard weighting approach in meta-analysis to manage group-size discrepancies and enhance estimator precision (Hedges and Olkin 2014).
@@
and the appropriate weighted standard error (NIST 2003) is:
@@
Each model response in our dataset corresponds to a 6-point Likert-scale judgment, formally an ordinal variable. Nonetheless, parametric methods such as the t-test can reliably analyze Likert-type data when averaged over moderate-to-large samples (Norman 2010; de Winter and Dodou 2010). In our experiments, each model–gender–claim cell aggregates at least 10 LLM completions, and statistics are computed at the level of topic means, averaged over individual claims, ensuring approximate normality via the Central Limit Theorem.
@@
  • Erosion on divisive topics. Other models—including GPT-4, LLaMa3, Titan, and Mistral—tend to erode known gender differences, even on topics with well-established disagreement. Quality Check 2a shows that GPT-4 and Mistral fail to reflect divergent opinions on abortion, and Quality Check 2b reveals that these models, in addition to Titan and LLaMa3, also underrepresent human gender-level disagreement on topics such as Black Americans, LGBTQ issues, and Illegal Immigration, yielding significant negative $\bar{g}_{\omega}$ scores ($p<\alpha$).
@@
In the previous section, we tested our quality checks using the backstory prompting approach across several different models. To assess alternative strategies, we benchmark this approach against two others prominent in the literature: fine-tuning (Namikoshi et al. 2024) and in-context learning (Karanjai et al. 2025). Specifically, we use only the GPT-3.5c checkpoint to control for variation across model architectures and focus on comparing simulation strategies. We test the following strategies:

### v4

@@ Abstract
The emergent capabilities of large language models (LLMs) have prompted interest in using them as surrogates for human subjects in opinion surveys. However, prior evaluations of LLM-based opinion simulation have relied heavily on costly, domain-specific survey data, and mixed empirical results leave their reliability in question. To enable cost-effective, early-stage evaluation, we introduce a quality control assessment designed to test the viability of LLM-simulated opinions on Likert-scale tasks without requiring large-scale human data for validation. This assessment comprises two key tests: logical consistency and alignment with stakeholder expectations, offering a low-cost, domain-adaptable validation tool. We apply our quality control assessment to an opinion simulation task relevant to AI-assisted content moderation and fact-checking workflows—a socially impactful use case—and evaluate nine LLMs using a baseline prompt engineering method (backstory prompting), as well as fine-tuning and in-context learning variants. None of the models or methods pass the full assessment, revealing several failure modes. We conclude with a discussion of the risk management implications and release TopicMisinfo, a benchmark dataset with paired human and LLM annotations simulated by various models and approaches, to support future research.
Code & Datahttps://tinyurl.com/qualitychecksAAAI
Full Paper with Appendiceshttps://arxiv.org/abs/2504.08954
@@ Introduction
Large language models (LLMs) have demonstrated significant appeal and versatility across a wide range of tasks. While the primary training objective of the base model is to predict the next most likely token, training on vast and diverse datasets, combined with increasingly complex system architectures, has resulted in the “emergence” of capabilities that were not explicitly anticipated during development (wei2022emergent). For instance, LLMs have been shown to learn new tasks through in-context learning (brown2020fewshot), perform coding tasks with remarkable accuracy (zheng2024survey), and have displayed human-like capabilities across different tasks (kosinski2024evaluating; strachan2024testing).
@@
These emerging capabilities have led researchers to investigate more unconventional uses of LLMs, including automating opinion surveys by simulating the viewpoints of specific demographic or ideological groups (qu2024performance). The possibility that LLMs may successfully simulate demographic viewpoints stems from (i) evidence that LLMs can be prompted to replicate responses from human subject studies in psychology and other social sciences (aher2023simulate), (ii) training data that likely encompass a broad spectrum of perspectives and opinions (miranda2024diversity), and (iii) prompting techniques that lead LLMs to produce seemingly coherent, opinionated responses (wright-etal-2024-llm). Automating such surveys could have broad applications in marketing (sarstedt2024marketing), content moderation (frohling2024personas), policymaking, and public relations (sanders2023demonstrations). LLMs may be especially valuable for pilot studies (sarstedt2024marketing; rothschild2024opportunities), sampling hard-to-reach populations (jansen2023employing), and labeling data that may be psychologically harmful for human annotators (wang2025large).
@@
However, existing evidence is inconclusive on the extent to which LLM-based methods can accurately simulate human opinions. Some studies show that including a demographic “backstory” in a prompt (Argyle2023simulate; bui2025mixture; jiang2024donald) (i.e., “backstory” prompting), fine‑tuning (namikoshi2024using), and in‑context learning (karanjai2025synthesizing) are approaches that can successfully approximate group‑level opinions, while others document systematic failures of these same approaches—especially when simulating the opinions of minority or non‑Western groups (santurkar2023aligning; sun2023aligningwhomlargelanguage; qu2024performance; mingmeng2024large; orlikowski2025beyond). Additional work finds that backstory prompting approaches—the most prevalent approach to simulating opinions in prior research—flatten within-group variance of opinions (wang2025large; Bisbee2024Synthetic; mingmeng2024large) and fails to generalize across topics (sanders2023demonstrations; lee2024can). See Appendix A for a literature review table that summarizes prior works methodologies, respective application domains, and findings.
@@
The mixed results in the empirical literature suggest that the reliability of LLM-simulated opinions cannot be assumed across domains and instead requires rigorous validation for each new use case. Importantly, however, current validation methods rely on gathering high-quality human-labeled data—the cost of which can rival that of traditional surveys. For early‑stage assessments, this upfront expense is likely to outweigh uncertain benefits, significantly hindering research and development efforts. The existence of applications with broad consensus about the potential desirability of simulating opinions, such as when human survey subjects would be exposed to harmful content (wang2025large) or trade secrets (wang2024large), motivates the need for validation techniques that lower the cost of early-stage exploration of LLM usage to simulate opinions.
@@
    Logical Consistency—whether models reliably position “average” opinions as a convex combination of group-level opinions and/or whether these average opinions are consistent with sampling from a stable reference population.
@@
The design of the proposed quality control assessment is grounded in the fact that some of the challenges involved in validating LLM-simulated opinions are similar to those involved in validating subjective human annotations. Researchers have developed various quality controls to assess the reliability of human data and annotator attentiveness (artstein2008inter; sap-etal-2022-annotators; davani-etal-2022-dealing; lease2011quality). Our proposed tests draw from this literature, adapting it to be suitable for assessing LLMs, with a focus on Likert-scale opinions. In the discussion section, we provide recommendations for integrating these tests into organizational risk management practices.
The value of these tests compared to those proposed in previous work (aher2023simulate; Argyle2023simulate; sun2023aligningwhomlargelanguage) is that they can be performed prior to expensive large-scale human validation effort, fostering well-informed deliberation on the suitability of LLM opinion simulation in a particular context.
@@
To illustrate the utility of our quality checks, we develop a domain-specific testbed in content moderation, focusing on the prioritization of potentially harmful misinformation for fact-checker review. Because prioritization is opinion-driven, resource-intensive, and exposes annotators to psychologically taxing content (liu2024exploring), it provides a compelling, socially impactful context for evaluating simulated opinions.
An effective testbed for group-level opinion simulation should include topics with both expected agreement and disagreement amongst groups. Gender offers strong prior expectations based on a significant body of public opinion research (huddy2008gender; lizotte2020gender). We focus on the prioritization criterion of identifying “potential harm to specific groups,” as previous research highlights this dimension of prioritization as one along which gender differences are particularly pronounced  (huddy2008gender; lizotte2020gender). [^1] While not the only criteria relevant for prioritization, research has noted that this is an important axis of consideration used by fact-checkers (sehat2024harm; liu2024exploring).
We built TopicMisinfo, a 160‑claim dataset spanning topics that can be expected to elicit gender disagreement and consensus. The selection of these topics is anchored in findings from the American National Election Studies (ANES) (ANES2021), which help us capture a diversity in topics that are likely to provoke varying degrees of gendered disagreement. We also included “Gold” comprehension checks, which are claims that are harmless, and are either obviously true (e.g., “A circle is round.”) or obviously false (e.g., “The tallest tree on Earth touches Mars.”), with a very strong prior on gender-level consensus of opinion. We also collect a small sample ($\approx 1600$ annotations) of task-specific human opinion data. Dataset composition, example claims, and human annotation statistics are detailed in the Appendices B-C; the TopicMisinfo dataset is publicly available to support future research.
@@
Applying the proposed quality checks, we evaluate simulated opinions from several commercially available LLMs: GPT-3.5-Turbo (checkpoints 06-13-23, 11-06-23, 01-25-24), GPT-4, GPT-4.1, GPT-5-mini, LLaMA 3, Titan-Text-Premier, and Mistral-Large—referred to throughout as GPT3.5a, GPT3.5b, GPT3.5c, GPT-4, GPT-4.1, GPT-5-mini, LLaMA3, Titan, and Mistral. We elicit group-level responses via backstory prompts (Argyle2023simulate; sun2023aligningwhomlargelanguage) that instruct models to adopt personas of men or women (see Appendix D for prompt templates). We also benchmark two other approaches to simulating opinions suggested in the literature: fine-tuning on small-sample human survey data (namikoshi2024using); and in-context learning with human-labeled examples (karanjai2025synthesizing). Importantly, our findings show that while fine-tuning and in-context learning generally improve alignment with stakeholder expectations, they do not substantially outperform backstory prompting on logical consistency. These results indicate that all evaluated models and prompting approaches fail the proposed quality control assessment for the task considered, suggesting that current methods face limitations in their ability to produce stable, internally coherent simulations of human opinions.
@@
In this section, we introduce two quality checks designed for early-stage evaluation of LLM-simulated opinions. For each check, we detail the methodology and analyze how a range of LLMs perform across gender conditions using the “backstory prompting” approach (Argyle2023simulate) on claims drawn from the TopicMisinfo dataset, focusing on a key opinion task relevant to fact-checking and content moderation workflows: assessing the potential harm of claims to specific groups. The code and reproducible pipeline are publicly available.
@@
Testing for logical consistency is a foundational tool for evaluating the quality of data from human annotators (aruguete2019serious) and from data aggregation methods such as majority voting (davani-etal-2022-dealing). In this work, we adapt this principle to assess the internal coherence of LLM-simulated opinions across demographic groups.
@@
For each claim $c_{i}$, the model is prompted $n_{i}$ times under each condition (men, women, average) using a temperature setting of $\tau=0.5$, generating distributions of Likert-scale responses. A logically consistent average opinion should be expressible as a convex combination of the group-level opinions it seeks to summarize. In the context of gender (assuming only two: men and women) [^2] While we assume a binary gender, Appendices F and G contain pseudo-code for generalizing this methodology for $\geq 2$ groups, while Appendix H includes the detail of a replication study including non-binary genders simulated with GPT3.5c., that would entail for each claim $c_{i}$, the model’s average predicted mean $\mu_{\text{avg}}(c_{i})$ should satisfy:
@@
We reject $H_{0}$ at level $\alpha$ if $\hat{p}(p_{0})<\alpha$, concluding that the model fails to produce average-conditioned opinions that lie within the convex hull of group-level distributions at an acceptable rate. We test acceptable rate values $p_{0}\in[0.7,0.8,0.9,1.0]$. See Appendix F for pseudocode that generalizes this test for $\geq 2$ groups.
@@
From this mixture, we resample synthetic average-conditioned responses, compute $\hat{q}^{(b)}_{c_{i}}$, and aggregate across claims: $\hat{L}^{(b)}_{\omega}(q_{0})=\frac{1}{N_{\omega}}\sum_{i=1}^{N_{\omega}}\left|\hat{q}_{c_{i}}^{(b)}-q_{0}\right|$. The $p$-value is $\Pr\left(\hat{L}^{(b)}_{\omega}(q_{0})>\hat{L}_{\omega}(q_{0})\right)$. This test is repeated for values of $q_{0}$ in $[0,1]$ at intervals of 0.05, applying a Bonferroni correction ($\alpha^{*}=0.0025$) to account for multiple comparisons. See Appendix G for pseudocode that generalizes this test for $\geq 2$ groups.
@@ Results.
Figure 1 reports feasible consistency thresholds ($p_{0}$) for the weak test and valid $q_{0}$ intervals for the strong test, indicating alignment with a consistent reference population. We highlight results for four socially divisive topics where group-level disagreement is well-documented. Given the lack of a gold standard for prompting average-conditioned or population-level perspectives, we evaluate consistency across four alternative prompts. See Appendix E for details.
  • GPT-3.5c: For threshold $p_{0}=0.7$, passes the weak test across three out of four topics with Base Prompt 4, indicating basic geometric plausibility. For the strong test, it only passes the Abortion topic, with feasible mixture weights $q_{0}\in[0,0.55]$ or $[0,0.85]$ depending on prompt. Fails to produce a consistent average-conditioned reference population for Black Americans, Illegal Immigration, and LGBTQ.
  • GPT-4.1: One prompt passes the weak test for Illegal Immigration, but we see no other feasible thresholds across the prompts and topics tested. For the strong test, we see the same prompt (Base Prompt 4) generate a plausible value for the topic of Illegal Immigration, but no other plausible values across topics and prompts.
  • GPT-5-mini: The only reasoning model we tested does not perform markedly better. One feasible point for the weak test for the LGBTQ topic, but no others. We see better performance on the strong test than other models, with feasible values for multiple prompts across two topics — Illegal Immigration and LGBTQ.
  • LLaMA-3: Two prompts pass the weak test at $p_{0}=0.7$ for Illegal Immigration and all four prompts pass for LGBTQ, but fails on other topics. For the strong test, exhibits moderate success: feasible $q_{0}\in[0,0.5]$ for Black Americans across prompts, and up-samples opinions of men for Illegal Immigration ($q_{0}\in[0.95,1]$). Fails on Abortion and LGBTQ.
@@
  • Titan: Two prompts pass the weak test at $p_{0}=0.7$ for Black Americans and one prompt passes for LGBTQ at $p_{0}=0.7$. On the strong test, Prompt 4 yields success on Black Americans with full-range plausibility ($q_{0}\in[0,1]$). Prompts 1 and 2 yield plausible $q_{0}\ \in[0,1]$ for LGBTQ, but it fails on both Abortion and Illegal Immigration.
@@
Assessing alignment with common sense or widespread expectations is standard for evaluating human annotations, e.g., via attention checks (lease2011quality). The same underlying idea can be leveraged when assessing the quality of LLM-generated annotations. In our setting, for instance, stakeholders—e.g., end-users or developers—can have clear intuitions about cases where group-level differences should or should not arise. Checking for misalignment on these cases can provide a powerful lens for evaluating the quality of LLM-generated annotations. If neglected, such misalignment can bias downstream tasks, leading to skewed decisions or misrepresentation of group perspectives. In high-stakes settings like claim prioritization for fact-checking, such errors carry serious ethical consequences.
@@
Topic gpt-3.5a gpt-3.5b gpt-3.5c gpt-4 gpt-4.1 gpt-5-mini llama3 mistral-large titan-text-premier
$\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val $\bar{g}_{\omega}$ p-val
@@
Abortion 1.65 * 0.28 0.60 -1.70 *** -1.34 *** -1.89 *** -1.65 *** -1.24 ** -1.16 *
Black Americans 1.72 * 0.35 0.39 -1.28 *** -1.01 *** -0.92 ** -1.09 *** -1.92 * -0.97 **
Illegal Immigration 0.92 ** -1.06 * -0.61 -1.42 ** -1.05 ** -1.04 ** -0.80 * -0.43 -0.71 *
LGBTQ 0.66 -1.37 * -0.37 -0.92 * -1.06 ** -0.99 * -1.08 ** -0.99 ** -1.10 **
@@
Entertainment 1.59 * 1.05 * 1.18 * 0.35 ** 0.39 ** 0.47 ** 0.41 ** 0.58 * 0.48 *
Gold 0.43 0.34 0.71 ** -0.01 0.10 0.10 -0.30 0.02 0.32
HealthScience 1.86 * 1.07 * 0.65 -0.19 0.58 * 0.32 0.04 0.79 0.56
Sports -0.01 -0.06 0.21 -0.26 -0.26 -0.36 -0.26 -0.26 -0.16
USA 1.68 *** 1.12 *** 0.63 ** -0.61 *** -0.05 -0.15 -0.21 -0.56 * 0.16
WeatherClimate 1.06 * 1.27 * 0.97 * -0.39 0.01 0.06 -0.10 -0.68 0.10
@@
Table 2: Quality Check 2b Statistical Results. GPT-3.5-turbo models tend to significantly exaggerate gender differences on topics humans found less divisive, while GPT-4, GPT-4.1, GPT-5-mini, Llama3, mistral-large, and Titan tended to significantly erode differences on topics humans found more divisive.
@@
a standard weighting approach in meta-analysis to manage group-size discrepancies and enhance estimator precision (hedges2014statistical).
@@
and the appropriate weighted standard error (nistweightsd) is:
@@
Each model response in our dataset corresponds to a 6-point Likert-scale judgment, formally an ordinal variable. Nonetheless, parametric methods such as the t-test can reliably analyze Likert-type data when averaged over moderate-to-large samples (norman2010likert; deWinter2010likert). In our experiments, each model–gender–claim cell aggregates at least 10 LLM completions, and statistics are computed at the level of topic means, averaged over individual claims, ensuring approximate normality via the Central Limit Theorem.
@@
  • Erosion on divisive topics. Other models—including GPT-4, GPT-4.1, GPT-5-mini, LLaMa3, Titan, and Mistral—tend to erode known gender differences, even on topics with well-established disagreement. Quality Check 2a shows that GPT-4, GPT-5-mini, and Mistral fail to reflect divergent opinions on abortion, and Quality Check 2b reveals that these models, in addition to GPT-4.1, Titan and LLaMa3, also underrepresent human gender-level disagreement on topics such as Black Americans, LGBTQ issues, and Illegal Immigration, yielding significant negative $\bar{g}_{\omega}$ scores ($p<\alpha$).
@@
In the previous section, we tested our quality checks using the backstory prompting approach across several different models. To assess alternative strategies, we benchmark this approach against two others prominent in the literature: fine-tuning (namikoshi2024using) and in-context learning (karanjai2025synthesizing). Specifically, we use only the GPT-3.5c checkpoint to control for variation across model architectures and focus on comparing simulation strategies. We test the following strategies: